what is the slope and y intercept of 5x=9y+81?
Answer by unlockmath(1688)

Hello,
First, let's set up the equation 5x=9y+81 in this form:
y=mx+b To do that subtract 9y from both sides:
5x-9y=81 Then subtract 5x from both sides:
-9y=-5x+81 Now divide both sides by -9 to give us:
y=5/9(x)-9 Ok, now we have it in the form as above.
From this we can determine the slope which is 5/9 and the y intercept is -9.
Make sense?
RJ
Check out a new math book at:
www.math-unlock.com